[figure: 9.7 cm]

Magnetic structures and phase transitions in TbRhSi, DyRhSi and

HoRhSi.

Ł. Gondek^{a}, B. Penc^{a}, N. Stüsser^{b}, A. Szytuła^{a} and A. Zygmunt^{c} aInstutute of Physics, Jagiellonian University, Reymonta 4, 30-059 Kraków

bBerlin Neutron Scattering Center, Hahn-Meitner Institute, D-14109 Berlin-Wannsee,

Germany

cW. Trzebiatowski Institute of Low Temperatures and Structure Research, Polish Academy of Sciences, 50-950 Wrocław

Neutron diffraction and magnetometric measurements show that RRhSi (R = Tb, Dy, Ho) are antiferromagnets below the Néel temperature equal 26 K for TbRhSi, 14.6 K for DyRhSi and 8.5 K for HoRhSi. At 1.5 K the magnetic structure of TbRhSi is described as a sine-wave modulated one with the propagation vector k⃗=(0;0.393(2);0.204(2)). Near by 16 K it transforms into a magnetic structure represented by the propagation vector k⃗=(0;0.3878(3);0). Simultanously a change of the direction of the magnetic moment is observed. For DyRhSi at T_{t} equal 5.4 K the magnetic structure changes from a collinear commensurate one, described by the propagation vector k⃗=(\frac{1}{2};0;\frac{1}{2}), to the incommensurate sine-wave structure described by the propagation vector k⃗=(0.418(2);0;0.515(2)). In both phases Dy magnetic moments are parallel to the b-axis. For HoRhSi magnetic structure similar to this observed for DyRhSi is stable in thr temperature range from 1.5 K up to T_{N}.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : A. Szytuła

Address for correspondence : Instutute of Physics, Jagiellonian University, Reymonta 4, 30-059 Kraków

Email address : szytula@if.uj.edu.pl